OpenStudy (anonymous):

Is the system consistent or inconsistent? 2x + y = 7 and x - 2y = 7

OpenStudy (anonymous):

consistent

OpenStudy (anonymous):

cute default

OpenStudy (anonymous):

this is consistent and independent x=21/5 y=-7/5